### v2

@@ Appendix H Centaur Experiments
We estimated a linear mixed-effects regression with participant as a random effect, log-transformed price perceptions as the dependent variable, and Participant Type (Human =0=0, Digital Twin =1=1), Digital Certification (Control =0=0, Digital Passport =1=1), and their interaction as predictors. We found a significant effect of Digital Certification ($b=.25$, $SE=.05$, $t(1190.45)=4.94$, $p<.001$), a significant effect of Participant Type ($b=.83$, $SE=.05$, $t(597.99)=16.91$, $p<.001$), and a significant interaction (b=0.21b=-0.21, $SE=.07$, t(596.99)=3.08t(596.99)=-3.08, $p<.01$; Figure 33), indicating that, at an aggregate level, the digital twins rated price perceptions differently than the humans. Pairwise comparisons confirmed that, for humans, price perceptions were significantly higher when their purchase was connected to a digital certification ($M=5.26$, $SE=.04$) compared to no digital certification ($M=5.01$, $SE=.04$; t(1190)=4.94t(1190)=-4.94, $p<.001$). However, for the digital twins, price perceptions were not significantly different when their purchase was connected to a digital certification ($M=5.88$, $SE=.04$) compared to no digital certification ($M=5.84$, $SE=.04$; t(1190)=0.73t(1190)=-0.73, $p=.463$).
@@
We next estimated a linear mixed-effects regression with participant as a random effect, status perceptions as the dependent variable, and Participant Type (Human =0=0, Digital Twin =1=1), Digital Certification (Control =0=0, Digital Passport =1=1), and their interaction as predictors. We found a marginally significant effect of Digital Certification ($b=0.14$, $SE=0.10$; $t(1175.69)=1.44$; $p=.151$), along with a significant effect of Participant Type ($b=0.19$, $SE=0.09$; $t(597.99)=2.06$; $p=.040$) and a significant interaction (b=0.55b=-0.55, $SE=0.13$; t(597.99)=4.24t(597.99)=-4.24; $p<.001$; Figure 34), indicating that, at an aggregate level, the digital twins rated status perceptions differently than the humans. Pairwise comparisons confirmed that, for humans, status perceptions were marginally higher when their purchase was connected to a digital certification ($M=4.81$, $SE=0.07$) compared to no digital certification ($M=4.66$, $SE=0.07$; t(1176)=1.44t(1176)=-1.44; $p=.150$). However, for the digital twins, status perceptions were significantly lower when their purchase was connected to a digital certification ($M=4.44$, $SE=0.07$) compared to no digital certification ($M=4.85$, $SE=0.07$; $t(1176)=4.16$; $p<.001$).
@@
These findings demonstrate that providing digital certification via a digital passport increases the value perception of a diffused luxury product for humans. Digital twins’ value perceptions, however, are not impacted by the inclusion of a digital passport. The findings have important implications for our understanding of when digital twins will not act as their human counterparts.
In this study, the divergence may stem from the psychological reassurance of product authenticity that certification provides to human consumers [42], which digital twins cannot fully internalize. This highlights a potential blind spot in current AI models: a limited ability to model symbolic consumption, ownership signaling, or status restoration processes. More broadly, these findings call for caution when relying on digital twins to simulate human judgments in domains involving identity and social perception—especially in luxury or prestige-based contexts. Future work should explore how these divergences arise and whether digital twins can be improved to better simulate symbolic inference processes, or whether some aspects of human value perception are fundamentally non-transferrable.
@@
Hidden fees and surcharges are a routine but often resented aspect of modern commerce. Whether booking a hotel, ordering food delivery, or using a credit card bill, consumers frequently encounter extra, often unexpected charges—like resort fees, service fees, foreign transaction fees, or order processing fees—added on top of the advertised base price. These fees are often disclosed late in the transaction, hindering consumers’ ability to estimate total costs or compare options, and are often described using obtuse language, obscuring understanding of why the fees are assessed. As a result, they have attracted growing scrutiny from consumer advocates and regulators. In recent years, a bipartisan group of regulators in the United States and other countries have advocated for stronger pricing transparency and restrictions on the delayed disclosure of mandatory charges. In 2022, the Biden-Harris administration labeled these “junk fees,” defined as “fees designed either to confuse or deceive consumers or to take advantage of lock-in or other forms of situational market power” [55].
@@
In the domain of policy attitudes, digital twins again tracked human performance. For both groups, conservatism was correlated with less support for pricing regulation, replicating prior work on ideological influences in policy attitudes (e.g., [38]). However, digital twins exhibited exaggerated ideological consistency with a stronger negative association between conservatism and regulation support than human participants. This likely reflects LLM’s reliance on dominant associations in their training data (e.g., conservatism = free-market values), whereas human responses may incorporate more contextual nuance.
@@
    • Question: "Based on the chapter you just read, what do you think is the likelihood that the text of the next chapter will be very negative or very positive? Please assign a percentage to each. The total must sum to 100%."
@@
    • Question: "Based on the chapter you just read, what do you think is the likelihood that the text of the next chapter will be very low energy or very high energy? Please assign a percentage to each. The total must sum to 100%."
@@
    • "How important is it to you that your employer actively invests in environmental sustainability (e.g., reducing carbon emissions, minimizing waste)?”
    • "Would you prefer to work at a company that is outspoken about social and political issues, or one that remains neutral?"
@@
    • "How likely are you to accept a job that requires you to be in the office five days a week?"
    • "Which of the following is more important to you when choosing a job? Where options included: (i) A flexible schedule that allows me to manage personal responsibilities; (ii) Clear boundaries between work and personal time, (iii) Strict 9-to-5 hours (iv) A predictable, fixed schedule [^20] To construct the index measuring respondents’ preferences for flexible work–life balance arrangements, we ordered the four response options from least to most flexible as follows: (1) Strict 9–5 hours, (2) Predictable, fixed schedule, (3) Clear boundaries between work and personal time, and (4) Flexible schedule that allows me to manage personal responsibilities. This index captures increasing preference for autonomy in managing one’s work schedule.
@@
    • "How much do you value opportunities for formal training and upskilling in your current or future role?"
    • "Imagine two job offers: one from a prestigious company with a competitive culture, and one from a lesser-known company with a collaborative, supportive environment. Which would you choose?" [^21] To construct the index for collaborative culture preference, we ordered the responses to reflect increasing preference for collaborative environments: (1) Prestigious, (2) Depends on compensation, (3) Not sure, and (4) Collaborative. This index captures respondents’ orientation toward workplace culture, with higher scores indicating stronger preferences for collaboration over prestige or ambiguity.
@@
    • "How important is it that your company’s leadership communicates transparently about the firm’s goals and challenges?"
    • "How much do you care about your company’s stated mission and values aligning with your personal beliefs?"
@@
Our results highlight several opportunities and drawbacks for the use of digital twins in human resource management. In terms of predicting preferences, we find that digital twins over-estimate human preferences for workplace attributes like collaboration, leadership transparency, and sustainability. Meanwhile, they are more accurate in estimating preferences for work-life balance, flexibility, and career development opportunities. This pattern suggests that while digital twins may be able to capture workplace preferences related to instrumental outcomes, they may struggle in domains with socially oriented or relational preferences, which are fundamental to modern-day workplaces. [^22] See, for example, [15], who show how automation interferes with workers’ preference for sociality in the workplace. Instead, digital twins over-estimate the importance of these attributes, relative to human jobseekers who are driven less by these attributes.
As both job candidates and employers adopt digital twins, understanding how digital twins shape the matching process is of utmost importance. Our results indicate digital twins can predict workplace preferences for job seekers in some domains, though future work remains. For example, one important future direction is understanding the training data necessary for digital twins to work at scale. While our setting features a broad array of background information used to program the digital twin [56], more targeted data collection, for example through the evaluation of hypothetical job postings, may increase the performance of digital twins in the hiring domain (see, for example, [33]). Another important avenue for future work is to understand how the labor market influences the performance of digital twins. For example, while digital twins may be able to predict preferences over short-term task assignments [50], they may struggle with more formal job opportunities, which may require a deeper appreciation of tradeoffs between financial incentives, meaning at work, pro-social tendencies, and other attributes.
@@
Second, research has shown that humans tend to show AI aversion or human favoritism when evaluating the creativity of ideas [65], i.e., giving higher creativity ratings to ideas coming from humans when informed that the ideas come from humans (as opposed to AI). We explore whether the reverse is true for digital twins, i.e., do digital twins show human favoritism, AI favoritism, or neither when evaluating ideas?
@@
Misinformation is a significant societal problem. Prior research has suggested that consumers share misinformation because they aren’t trained [37], predisposed [44], prompted [45], or incentivized [12] to consider accuracy in online settings. Accuracy, however, is only one consideration people may have when sharing news with their online social network. Recent research suggests that people may prioritize other goals (e.g., entertainment) over accuracy when sharing news online, even when explicitly prompted with accuracy cues [34]. The current research aims to replicate this pattern with human participants, and test whether their digital twins make a similar trade-off. Specifically, this study tests how people prioritize five important considerations when sharing article headlines using a ratings-based conjoint design: headline entertainingness, source trustworthiness, article content, political lean, and number of likes.
These attributes were selected to capture a range of factors known to influence online content sharing decisions. Headline entertainingness (more vs. less entertaining) was included to reflect the growing role of emotional engagement in news virality [3] and the spread of misinformation online [58]. Source trustworthiness (trustworthy vs. untrustworthy) was included to test whether individuals prioritize credibility when deciding what to share (e.g., [43]. Content type (entertaining vs. informative) was included to disentangle preferences for how the headline is written (e.g., more or less entertaining) from article content. Political lean (conservative vs. liberal) was included to account for the influence of ideological alignment and partisan identity on engagement with news and misinformation [57]. Finally, the number of likes (20 vs. 200 vs. 2000) served as a proxy for popularity and social proof [12].
@@
After the sharing task, participants stated their perceptions of each attribute in terms of importance and accuracy diagnosticity. For each of the attributes, participants were asked, “How important are each of the following attributes when you are considering sharing news articles with other people on social media?” (1-Not important at all, 7-Extremely important). Participants were then asked, “To what extent can you use each of these attributes to determine whether a news article is accurate?” (1-Not at all, 7-Extremely). To test the extent to which participants make a conscious decision to prioritize either accuracy or entertainment, they were additionally asked, “When you share news with others on social media, is it more important that the content is:” (1-Verifiably correct, 6-Entertaining; [34]).
@@
Part-worth utilities for each attribute were calculated using Ordinary Least Squares (OLS) regression, a widely used approach in marketing research. Specifically, individual-level part-worth utilities were estimated for each attribute level using linear regression, treating the likelihood rating as the dependent variable and dummy-coded attribute levels as predictors. These regressions were run separately for each individual (human or twin). In a subsequent step, we computed the attribute range for each individual (maximum – minimum utility within each attribute) and then ran separate OLS regressions to estimate the effect of Participant Type (human vs. twin) on each attribute’s range, yielding the “Twin–Human $\Delta$” along with associated standard errors, confidence intervals, and $p$-values. The approach follows recommendations from [41].
@@
All pairwise comparisons are reported in Table LABEL:SI_infotainment_tab3. For humans, more entertaining headlines were marginally preferred more over less entertaining headlines. Human participants also significantly preferred entertaining over informative content and conservative over liberal content. Preference for trustworthy over untrustworthy sources was not statistically significant. Finally, a non-linear pattern emerged for number of likes: humans showed a significant preference for 2,000 likes over 200 likes, but non-significant preference for 2,000 over 20 likes and 200 over 20 likes. Twins, on the other hand, showed a significant preference for less entertaining headlines and significant preference for informative over entertaining content, but in contrast to humans, held a strong and significant preference for trustworthy over untrustworthy sources and liberal over conservative content. Twins also showed striking and consistent preferences for higher like counts across all comparisons, particularly between 20 and 2000 likes (99.33%), highlighting a stronger responsiveness to this popularity signal than humans.
@@
The above analyses assess participants’ perceptions of attribute importance and accuracy diagnosticity using continuous scales that do not force trade-offs. To test the extent to which participants believe accuracy or entertainment is relatively more important when sharing articles online, a repeated-measures ANOVA of participant type (human vs. twin) on the relative preference question was next conducted: “When you share news with others on social media, is it more important that the content is:” (1 = Verifiably correct, 6 = Entertaining; working paper, Lane and Brucks, 2025). The results revealed that humans ($M=2.83$, $SD=1.73$) were significantly more likely than their digital twins ($M=2.12$, $SD=0.61$) to report that accuracy is more important than entertainment when sharing news on social media, $t(299)=7.01$, $p<.001$, $d=0.40$. This suggests that humans believe they place greater emphasis on factual correctness, whereas twins show comparatively more actual emphasis on factual correctness in the conjoint analysis.
@@
These findings are surprising in light of prior literature that suggests highlighting accuracy considerations can curb the sharing of low quality or false news (e.g., [43]). When balancing both accuracy and entertainingness considerations, however, it seems that people prioritize headline entertainingness over accuracy (i.e., source). Thus, the findings suggest that accuracy nudges may not overwhelm a desire to share entertaining news, replicating recent research (working paper, [34]). Twins, in contrast, behave more like what prior literature would expect, balancing headline entertainingness with positive social feedback [12] and prioritizing trustworthy over untrustworthy sources (e.g.,[43]). Finally, twins assume their human counterparts are more liberal leaning than they actually are, supporting earlier findings demonstrating twins to have more liberal views (e.g., pro-vaccination, pro-immigration) compared to their underlying humans [56].
Divergence between humans and twins may arise from how predictive models infer behavior. Digital twins are built to generalize from past behavior and stated attitudes, so they assume consistency across different informational contexts. As a result, they may impose a kind of rationality on human behavior that reflects canonical cognitive psychology theories: weighting political alignment and popularity as key motivators and distinguishing source trustworthiness when it is made salient. Humans, on the other hand, seem to display more variance in their decisions, which may be guided by momentary salience or idiosyncratic preferences, including both personal preferences and varying social preferences or pressures. For example, the effect of all article attributes is flattened among twins, compared to humans. This smoothing may reflect the averaging tendency of algorithmic inference, whereas humans rely on heuristics that elevate certain features above others depending on the specific moment, task, or social considerations such as their audience [34]. Ultimately, this work suggests that human preferences for content sharing are not only more nuanced than previously thought, but they are also harder to predict, even by a system trained on the humans themselves.
@@
There is a growing body of literature on the creative capabilities of LLMs. This literature typically shows that LLMs are able to generate ideas that are judged as creative or more creative than ideas generated by humans (e.g., [7, 35, 65]). The first question this study seeks to address is: How do digital twins’ creative abilities compare to those of their human counterparts?
@@
    Divergent Association Task (DAT, Olson et al., 2021), which measures divergent thinking abilities (single task - "Please write 10 words that are as different from each other as possible, in all meanings and uses of the words."),
@@
    Shortest Semantic Path Task (SSPT, Toubia and Berger, 2025), which may be viewed as a measure of convergent thinking abilities (5 tasks, using the following pairs of seed words: eternity-curiosity, elephant-galaxy, perseverance-eloquence, euphoria-tulip, tangerine-penguin: "Find a way to connect these two words: <seed1> and <seed2>. Each word in the sequence below should be as closely related as possible to the word before it. The first word in the sequence is already set to <seed1>. The last word is already set to <seed2>. That is, you only need to add 3 words that connect the first word to the last word."),
@@
    An idea generation task asking for a new idea (at least 200 characters) for smartphone apps that will help their users keep a healthier lifestyle (single task: "HOW COULD SMARTPHONES HELP THEIR USERS BE HEALTHIER? We are interested in new ideas for smartphone apps that will help their users keep a healthier lifestyle. In the space provided below, please enter one new idea for a smartphone app priced at \$0.99 that would help its users keep a healthier lifestyle. Please be as specific as possible and describe the main features of the app. Please enter exactly one detailed idea in the box below. Your idea should contain at least 200 characters. ").
@@
In addition, this study leverages the Forward Flow measure [20] collected in the original Twin2K500 dataset. Performance on this task is measured as the average pairwise semantic distance between the words in the sequence.
@@
    Self-persuasion [8]. When prompted to consider the other side, do digital twins abandon their attitudes more readily than their human counterparts?
@@
For the first task, we followed the self-persuasion condition in [10]. Specifically, all participants first indicated their attitude toward universal basic income (UBI) using a 0-100 scale in response to the statement “America should have a universal basic income system”? where 0 = Strongly disagree (I am against universal basic income) and 100 = Strongly agree (I am in favor of universal basic income). Then, following the attitude measure, participants were instructed to take the alternate point of view. Due to technical constraints with dynamic surveys, the wording differed slightly between human and digital twin participants.
@@
We also examined the language used when considering the other side. Specifically, we measured the language similarity between a human participant’s text response and their twin’s response versus the text response of a randomly selected twin using the BERT metric, defined as “the sum of cosine similarities between their tokens’ embeddings” [64] In a random-effects regression with language matching as the dependent variable; same-twin versus different-twin pairing as the independent variable; and human participant identifier as a random effect, we found that the language between a human and their twin was slightly more similar ($M=0.19$, $SD=0.09$) than between a human and a random twin ($M=0.18$, $SD=0.09$; $b=0.009$, $SE=0.001$; $t(1000)=5.84$; $p<.001$).

### v4

@@ Appendix H Centaur Experiments
We estimated a linear mixed-effects regression with participant as a random effect, log-transformed price perceptions as the dependent variable, and Participant Type (Human =0=0, Digital Twin =1=1), Digital Certification (Control =0=0, Digital Passport =1=1), and their interaction as predictors. We found a significant effect of Digital Certification ($b=.25$, $SE=.05$, $t(1190.45)=4.94$, $p<.001$), a significant effect of Participant Type ($b=.83$, $SE=.05$, $t(597.99)=16.91$, $p<.001$), and a significant interaction (b=0.21b=-0.21, $SE=.07$, t(596.99)=3.08t(596.99)=-3.08, $p<.01$; Figure 33), indicating that, at an aggregate level, the digital twins rated price perceptions differently than the humans. Pairwise comparisons confirmed that, for humans, price perceptions were significantly higher when their purchase was connected to a digital certification ($M=5.26$, $SE=.04$) compared to no digital certification ($M=5.01$, $SE=.04$; $t(1190)=4.94$, $p<.001$). However, for the digital twins, price perceptions were not significantly different when their purchase was connected to a digital certification ($M=5.88$, $SE=.04$) compared to no digital certification ($M=5.84$, $SE=.04$; $t(1190)=0.73$, $p=.463$).
@@
We next estimated a linear mixed-effects regression with participant as a random effect, status perceptions as the dependent variable, and Participant Type (Human =0=0, Digital Twin =1=1), Digital Certification (Control =0=0, Digital Passport =1=1), and their interaction as predictors. We found a non-significant effect of Digital Certification ($b=0.14$, $SE=0.10$; $t(1175.69)=1.44$; $p=.151$), along with a significant effect of Participant Type ($b=0.19$, $SE=0.09$; $t(597.99)=2.06$; $p=.040$) and a significant interaction (b=0.55b=-0.55, $SE=0.13$; t(597.99)=4.24t(597.99)=-4.24; $p<.001$; Figure 34), indicating that, at an aggregate level, the digital twins rated status perceptions differently than the humans. Pairwise comparisons confirmed that, for humans, status perceptions were marginally higher when their purchase was connected to a digital certification ($M=4.81$, $SE=0.07$) compared to no digital certification ($M=4.66$, $SE=0.07$; $t(1176)=1.44$; $p=.150$). However, for the digital twins, status perceptions were significantly lower when their purchase was connected to a digital certification ($M=4.44$, $SE=0.07$) compared to no digital certification ($M=4.85$, $SE=0.07$; t(1176)=4.16t(1176)=-4.16; $p<.001$).
@@
These findings demonstrate that providing digital certification via a digital passport increases the value perception of a diffused luxury product for humans, consistent with existing research (?). Digital twins’ value perceptions, however, are not impacted by the inclusion of a digital passport. The findings have important implications for our understanding of when digital twins will not act as their human counterparts.
In this study, the divergence may stem from the psychological reassurance of product authenticity that certification provides to human consumers (?), which digital twins cannot fully internalize. This highlights a potential blind spot in current AI models: a limited ability to model symbolic consumption, ownership signaling, or status restoration processes. More broadly, these findings call for caution when relying on digital twins to simulate human judgments in domains involving identity and social perception—especially in luxury or prestige-based contexts. Future work should explore how these divergences arise and whether digital twins can be improved to better simulate symbolic inference processes, or whether some aspects of human value perception are fundamentally non-transferrable.
@@
Hidden fees and surcharges are a routine but often resented aspect of modern commerce. Whether booking a hotel, ordering food delivery, or using a credit card bill, consumers frequently encounter extra, often unexpected charges—like resort fees, service fees, foreign transaction fees, or order processing fees—added on top of the advertised base price. These fees are often disclosed late in the transaction, hindering consumers’ ability to estimate total costs or compare options, and are often described using obtuse language, obscuring understanding of why the fees are assessed. As a result, they have attracted growing scrutiny from consumer advocates and regulators. In recent years, a bipartisan group of regulators in the United States and other countries have advocated for stronger pricing transparency and restrictions on the delayed disclosure of mandatory charges. In 2022, the Biden-Harris administration labeled these “junk fees,” defined as “fees designed either to confuse or deceive consumers or to take advantage of lock-in or other forms of situational market power” (?).
@@
In the domain of policy attitudes, digital twins again tracked human performance. For both groups, conservatism was correlated with less support for pricing regulation, replicating prior work on ideological influences in policy attitudes (e.g., (?)). However, digital twins exhibited exaggerated ideological consistency with a stronger negative association between conservatism and regulation support than human participants. This likely reflects LLM’s reliance on dominant associations in their training data (e.g., conservatism = free-market values), whereas human responses may incorporate more contextual nuance.
@@
    • Question: ”Based on the chapter you just read, what do you think is the likelihood that the text of the next chapter will be very negative or very positive? Please assign a percentage to each. The total must sum to 100%.”
@@
    • Question: ”Based on the chapter you just read, what do you think is the likelihood that the text of the next chapter will be very low energy or very high energy? Please assign a percentage to each. The total must sum to 100%.”
@@
    • ”How important is it to you that your employer actively invests in environmental sustainability (e.g., reducing carbon emissions, minimizing waste)?”
    • ”Would you prefer to work at a company that is outspoken about social and political issues, or one that remains neutral?”
@@
    • ”How likely are you to accept a job that requires you to be in the office five days a week?”
    • ”Which of the following is more important to you when choosing a job? Where options included: (i) A flexible schedule that allows me to manage personal responsibilities; (ii) Clear boundaries between work and personal time, (iii) Strict 9-to-5 hours (iv) A predictable, fixed schedule [^19] To construct the index measuring respondents’ preferences for flexible work–life balance arrangements, we ordered the four response options from least to most flexible as follows: (1) Strict 9–5 hours, (2) Predictable, fixed schedule, (3) Clear boundaries between work and personal time, and (4) Flexible schedule that allows me to manage personal responsibilities. This index captures increasing preference for autonomy in managing one’s work schedule.
@@
    • ”How much do you value opportunities for formal training and upskilling in your current or future role?”
    • ”Imagine two job offers: one from a prestigious company with a competitive culture, and one from a lesser-known company with a collaborative, supportive environment. Which would you choose?” [^20] To construct the index for collaborative culture preference, we ordered the responses to reflect increasing preference for collaborative environments: (1) Prestigious, (2) Depends on compensation, (3) Not sure, and (4) Collaborative. This index captures respondents’ orientation toward workplace culture, with higher scores indicating stronger preferences for collaboration over prestige or ambiguity.
@@
    • ”How important is it that your company’s leadership communicates transparently about the firm’s goals and challenges?”
    • ”How much do you care about your company’s stated mission and values aligning with your personal beliefs?”
@@
Our results highlight several opportunities and drawbacks for the use of digital twins in human resource management. In terms of predicting preferences, we find that digital twins over-estimate human preferences for workplace attributes like collaboration, leadership transparency, and sustainability. Meanwhile, they are more accurate in estimating preferences for work-life balance, flexibility, and career development opportunities. This pattern suggests that while digital twins may be able to capture workplace preferences related to instrumental outcomes, they may struggle in domains with socially oriented or relational preferences, which are fundamental to modern-day workplaces. [^21] See, for example, (?), who show how automation interferes with workers’ preference for sociality in the workplace. Instead, digital twins over-estimate the importance of these attributes, relative to human jobseekers who are driven less by these attributes.
As both job candidates and employers adopt digital twins, understanding how digital twins shape the matching process is of utmost importance. Our results indicate digital twins can predict workplace preferences for job seekers in some domains, though future work remains. For example, one important future direction is understanding the training data necessary for digital twins to work at scale. While our setting features a broad array of background information used to program the digital twin (?), more targeted data collection, for example through the evaluation of hypothetical job postings, may increase the performance of digital twins in the hiring domain (see, for example, (?)). Another important avenue for future work is to understand how the labor market influences the performance of digital twins. For example, while digital twins may be able to predict preferences over short-term task assignments (?), they may struggle with more formal job opportunities, which may require a deeper appreciation of tradeoffs between financial incentives, meaning at work, pro-social tendencies, and other attributes.
@@
Second, research has shown that humans tend to show AI aversion or human favoritism when evaluating the creativity of ideas (?), i.e., giving higher creativity ratings to ideas coming from humans when informed that the ideas come from humans (as opposed to AI). We explore whether the reverse is true for digital twins, i.e., do digital twins show human favoritism, AI favoritism, or neither when evaluating ideas?
@@
Misinformation is a significant societal problem. Prior research has suggested that consumers share misinformation because they aren’t trained (?), predisposed (?), prompted (?), or incentivized (?) to consider accuracy in online settings. Accuracy, however, is only one consideration people may have when sharing news with their online social network. Recent research suggests that people may prioritize other goals (e.g., entertainment) over accuracy when sharing news online, even when explicitly prompted with accuracy cues (?). The current research aims to replicate this pattern with human participants, and test whether their digital twins make a similar trade-off. Specifically, this study tests how people prioritize five important considerations when sharing article headlines using a ratings-based conjoint design: headline entertainingness, source trustworthiness, article content, political lean, and number of likes.
These attributes were selected to capture a range of factors known to influence online content sharing decisions. Headline entertainingness (more vs. less entertaining) was included to reflect the growing role of emotional engagement in news virality (?) and the spread of misinformation online (?). Source trustworthiness (trustworthy vs. untrustworthy) was included to test whether individuals prioritize credibility when deciding what to share (e.g., (?)). Content type (entertaining vs. informative) was included to disentangle preferences for how the headline is written (e.g., more or less entertaining) from article content. Political lean (conservative vs. liberal) was included to account for the influence of ideological alignment and partisan identity on engagement with news and misinformation (?). Finally, the number of likes (20 vs. 200 vs. 2000) served as a proxy for popularity and social proof (?).
@@
After the sharing task, participants stated their perceptions of each attribute in terms of importance and accuracy diagnosticity. For each of the attributes, participants were asked, “How important are each of the following attributes when you are considering sharing news articles with other people on social media?” (1-Not important at all, 7-Extremely important). Participants were then asked, “To what extent can you use each of these attributes to determine whether a news article is accurate?” (1-Not at all, 7-Extremely). To test the extent to which participants make a conscious decision to prioritize either accuracy or entertainment, they were additionally asked, “When you share news with others on social media, is it more important that the content is:” (1-Verifiably correct, 6-Entertaining; (?)).
@@
Part-worth utilities for each attribute were calculated using Ordinary Least Squares (OLS) regression, a widely used approach in marketing research. Specifically, individual-level part-worth utilities were estimated for each attribute level using linear regression, treating the likelihood rating as the dependent variable and dummy-coded attribute levels as predictors. These regressions were run separately for each individual (human or twin). In a subsequent step, we computed the attribute range for each individual (maximum – minimum utility within each attribute) and then ran separate OLS regressions to estimate the effect of Participant Type (human vs. twin) on each attribute’s range, yielding the “Twin–Human $\Delta$” along with associated standard errors, confidence intervals, and $p$-values. The approach follows recommendations from (?).
@@
All pairwise comparisons are reported in Table LABEL:SI_infotainment_tab3. For humans, more entertaining headlines were marginally preferred more over less entertaining headlines. Human participants also significantly preferred entertaining over informative content and conservative over liberal content. Preference for trustworthy over untrustworthy sources was not statistically significant. Finally, a non-linear pattern emerged for number of likes: humans showed a significant preference for 2,000 likes over 200 likes, but non-significant preference for 2,000 over 20 likes and 200 over 20 likes. Twins likewise showed a significant preference for more entertaining headlines, but in contrast to humans, held a significant preference for informative over entertaining content, trustworthy over untrustworthy sources, and liberal over conservative content. Twins also showed striking and consistent preferences for higher like counts across all comparisons, particularly between 20 and 2000 likes (99.33%), highlighting a stronger responsiveness to this popularity signal than humans.
@@
The above analyses assess participants’ perceptions of attribute importance and accuracy diagnosticity using continuous scales that do not force trade-offs. To test the extent to which participants believe accuracy or entertainment is relatively more important when sharing articles online, a repeated-measures ANOVA of participant type (human vs. twin) on the relative preference question was next conducted: “When you share news with others on social media, is it more important that the content is:” (1 = Verifiably correct, 6 = Entertaining; (?)). The results revealed that humans ($M=2.83$, $SD=1.73$) were significantly more likely than their digital twins ($M=2.12$, $SD=0.61$) to report that accuracy is more important than entertainment when sharing news on social media, $t(299)=7.01$, $p<.001$, $d=0.40$. This suggests that humans believe they place greater emphasis on factual correctness, whereas twins show comparatively more actual emphasis on factual correctness in the conjoint analysis.
@@
These findings are surprising in light of prior literature that suggests highlighting accuracy considerations can curb the sharing of low quality or false news (e.g., (?)). When balancing both accuracy and entertainingness considerations, however, it seems that people prioritize headline entertainingness over accuracy (i.e., source). Thus, the findings suggest that accuracy nudges may not overwhelm a desire to share entertaining news, replicating recent research (working paper, (?)). Twins, in contrast, behave more like what prior literature would expect, balancing headline entertainingness with positive social feedback (?) and prioritizing trustworthy over untrustworthy sources (e.g., (?)). Finally, twins assume their human counterparts are more liberal leaning than they actually are, supporting earlier findings demonstrating twins to have more liberal views (e.g., pro-vaccination, pro-immigration) compared to their underlying humans (?).
Divergence between humans and twins may arise from how predictive models infer behavior. Digital twins are built to generalize from past behavior and stated attitudes, so they assume consistency across different informational contexts. As a result, they may impose a kind of rationality on human behavior that reflects canonical cognitive psychology theories: weighting political alignment and popularity as key motivators and distinguishing source trustworthiness when it is made salient. Humans, on the other hand, seem to display more variance in their decisions, which may be guided by momentary salience or idiosyncratic preferences, including both personal preferences and varying social preferences or pressures. For example, the effect of all article attributes is flattened among twins, compared to humans. This smoothing may reflect the averaging tendency of algorithmic inference, whereas humans rely on heuristics that elevate certain features above others depending on the specific moment, task, or social considerations such as their audience (?). Ultimately, this work suggests that human preferences for content sharing are not only more nuanced than previously thought, but they are also harder to predict, even by a system trained on the humans themselves.
@@
There is a growing body of literature on the creative capabilities of LLMs. This literature typically shows that LLMs are able to generate ideas that are judged as creative or more creative than ideas generated by humans (e.g., (?, ?, ?)). The first question this study seeks to address is: How do digital twins’ creative abilities compare to those of their human counterparts?
@@
    Divergent Association Task (DAT, Olson et al., 2021), which measures divergent thinking abilities (single task - ”Please write 10 words that are as different from each other as possible, in all meanings and uses of the words.”),
@@
    Shortest Semantic Path Task (SSPT, Toubia and Berger, 2025), which may be viewed as a measure of convergent thinking abilities (5 tasks, using the following pairs of seed words: eternity-curiosity, elephant-galaxy, perseverance-eloquence, euphoria-tulip, tangerine-penguin: ”Find a way to connect these two words: <seed1> and <seed2>. Each word in the sequence below should be as closely related as possible to the word before it. The first word in the sequence is already set to <seed1>. The last word is already set to <seed2>. That is, you only need to add 3 words that connect the first word to the last word.”),
@@
    An idea generation task asking for a new idea (at least 200 characters) for smartphone apps that will help their users keep a healthier lifestyle (single task: ”HOW COULD SMARTPHONES HELP THEIR USERS BE HEALTHIER? We are interested in new ideas for smartphone apps that will help their users keep a healthier lifestyle. In the space provided below, please enter one new idea for a smartphone app priced at \$0.99 that would help its users keep a healthier lifestyle. Please be as specific as possible and describe the main features of the app. Please enter exactly one detailed idea in the box below. Your idea should contain at least 200 characters. ”).
@@
In addition, this study leverages the Forward Flow measure (?) collected in the original Twin2K500 dataset. Performance on this task is measured as the average pairwise semantic distance between the words in the sequence.
@@
    Self-persuasion (?). When prompted to consider the other side, do digital twins abandon their attitudes more readily than their human counterparts?
@@
For the first task, we followed the self-persuasion condition in (?). Specifically, all participants first indicated their attitude toward universal basic income (UBI) using a 0-100 scale in response to the statement “America should have a universal basic income system”? where 0 = Strongly disagree (I am against universal basic income) and 100 = Strongly agree (I am in favor of universal basic income). Then, following the attitude measure, participants were instructed to take the alternate point of view. Due to technical constraints with dynamic surveys, the wording differed slightly between human and digital twin participants.
@@
We also examined the language used when considering the other side. Specifically, we measured the language similarity between a human participant’s text response and their twin’s response versus the text response of a randomly selected twin using the BERT metric, defined as “the sum of cosine similarities between their tokens’ embeddings” (?) In a random-effects regression with language matching as the dependent variable; same-twin versus different-twin pairing as the independent variable; and human participant identifier as a random effect, we found that the language between a human and their twin was slightly more similar ($M=0.19$, $SD=0.09$) than between a human and a random twin ($M=0.18$, $SD=0.09$; $b=0.009$, $SE=0.001$; $t(1000)=5.84$; $p<.001$).